# Benchmarking LLM Faithfulness in RAG with Evolving Leaderboards

### v1

@@ Abstract
Hallucinations remain a persistent challenge for LLMs. RAG aims to reduce hallucinations by grounding responses in contexts. However, even when provided context, LLMs still frequently introduce unsupported information or contradictions. This paper presents our efforts to measure LLM hallucinations with a focus on summarization tasks, assessing how often various LLMs introduce hallucinations when summarizing documents. We discuss Vectara’s existing LLM hallucination leaderboard, based on the Hughes Hallucination Evaluation Model (HHEM). While HHEM and Vectara’s Hallucination Leaderboard have garnered great research interest, we examine challenges faced by HHEM and current hallucination detection methods by analyzing the effectiveness of these methods on existing hallucination datasets. To address these limitations, we propose FaithJudge, an LLM-as-a-judge approach guided by few-shot human hallucination annotations, which substantially improves automated LLM hallucination evaluation over current methods. We introduce an enhanced hallucination leaderboard centered on FaithJudge  [^1] https://github.com/vectara/FaithJudge, alongside our current hallucination leaderboard [^2] https://github.com/vectara/hallucination-leaderboard, enabling more reliable benchmarking of LLMs for hallucinations in RAG.
@@
Manveer Singh Tamber1∗, Forrest Sheng Bao2, Chenyu Xu2,3, Ge Luo2, Suleman Kazi2, Minseok Bae4∗, Miaoran Li3∗, Ofer Mendelevitch2, Renyi Qu2, Jimmy Lin1 1 University of Waterloo  2 Vectara  3 Iowa State University  4 Stanford University Correspondence: mtamber@uwaterloo.ca, {forrest, suleman}@vectara.com
11footnotetext: Work done while at Vectara
@@ 1 Introduction
LLMs excel in various tasks, but frequently produce hallucinations, generating false or misleading information unsupported by provided contexts or world knowledge Ji et al. (2023); Huang et al. (2025); Lin et al. (2022); Tang et al. (2023). While Retrieval-Augmented Generation (RAG) approaches Guu et al. (2020); Lewis et al. (2020b); Shuster et al. (2021) attempt to mitigate hallucinations by grounding responses in external trusted contexts, they do not fully eliminate hallucinations, as LLMs often introduce details unsupported by retrieved contexts, misrepresent information, or generate outright contradictions Niu et al. (2024).
An ongoing challenge within RAG is ensuring context-faithfulness Niu et al. (2024); Jia et al. (2023); Ming et al. (2024). Detecting when LLMs deviate from the information in the provided context remains a difficult problem. Although there has been progress, hallucination detection methods, including fine-tuned detectors largely for evaluating summaries Zhou et al. (2021); Gekhman et al. (2023); Honovich et al. (2022); Zha et al. (2023); Tang et al. (2024a) and LLM-as-a-judge techniques Zheng et al. (2023); Luo et al. (2023); Jacovi et al. (2025), continue to struggle with the accurate identification of LLM-generated hallucinations.
In this paper, we study and aim to improve hallucination evaluation in RAG by building on prior work in summary consistency evaluation. We analyze the capabilities and limitations of current hallucination detection methods, including fine-tuned models such as Vectara’s Hughes Hallucination Evaluation Model (HHEM) Bao et al. (2024) and zero-shot methods using LLM judges.
To overcome the challenges of LLM-as-a-judge techniques for zero-shot hallucination detection and fine-tuned hallucination detection models, we introduce FaithJudge, an LLM-as-a-judge approach guided by few-shot human annotations of hallucinations. FaithJudge leverages labelled hallucinations from diverse LLM generations to automate the evaluation of LLMs on their propensity to hallucinate when summarizing the same articles or using the same articles to respond to queries. This approach results in notably higher agreement with human judgments compared with existing automated methods. Additionally, we introduce an enhanced hallucination leaderboard based on FaithJudge, enabling more reliable benchmarking of hallucinations in LLM-generated summaries and responses.
We discuss both Vectara’s existing hallucination leaderboard Hughes and Bae (2023) based on HHEM and our new leaderboard based on FaithJudge. Hallucinations within RAG remain frequent and problematic, even in leading LLMs. Our approach contributes toward more accurate hallucination evaluation, aiding the development of more trustworthy generative AI systems.
Vectara serves customers across diverse industries. For many of these customers, addressing hallucinations in LLM outputs is a critical priority. Driven by this customer-centric need, we developed our hallucination leaderboard and benchmarking methods, which are presented in this paper.
@@ 2 Background
Accurate hallucination detection is essential for reliably quantifying hallucination rates in LLMs. Numerous datasets have been developed for evaluating hallucinations in summarization tasks. Earlier datasets, such as SummaC Laban et al. (2022) and AggreFact Tang et al. (2023), aggregated multiple resources, standardized labels, and classification taxonomies. However, these primarily focused on summaries from pre-ChatGPT models like fine-tuned T5 Raffel et al. (2020), BART Lewis et al. (2020a), and PEGASUS Zhang et al. (2020) models, potentially limiting their relevance to contemporary LLMs that may produce more nuanced and difficult to identify hallucinations.
@@
Recent benchmarks address this limitation by incorporating summaries generated by modern LLMs. TofuEval Tang et al. (2024b) provided hallucination labels on topic-focused dialogue summarization tasks with LLMs including GPT-3.5-Turbo, Vicuna Chiang et al. (2023) and WizardLM Xu et al. (2023). Similarly, HaluEval Li et al. (2023) included ChatGPT-generated hallucinations across summarization, question answering (QA), and dialogue tasks, while RAGTruth Niu et al. (2024) also annotated responses from models including GPT-3.5, GPT-4 OpenAI (2023), Llama-2 Touvron et al. (2023), and Mistral Jiang et al. (2023). FaithBench Bao et al. (2025) presents human annotations of challenging hallucinations in summaries from 10 modern LLMs from 8 different model families (detailed further in Section 4).
@@
Due to limited large-scale, human-annotated data for training hallucination detectors, early detection methods relied heavily on natural language inference (NLI) or question-answering (QA) systems Fabbri et al. (2022). For instance, SummaC aggregated sentence-level NLI entailment scores between document-summary sentence pairs. AlignScore Zha et al. (2023) extended this by training detection models on multiple semantic alignment tasks evaluated at the chunk level. MiniCheck Tang et al. (2024a) addressed data scarcity by synthesizing hallucinated examples using GPT-4 for model training.
@@
Modern LLMs’ strong zero-shot instruction-following capabilities have also enabled LLM-as-a-judge methods Zheng et al. (2023); Luo et al. (2023); Jacovi et al. (2025); Gao et al. (2023). Instead of evaluating entire generated summaries, approaches like FACTSCORE Min et al. (2023) and RAGAS Es et al. (2024) decompose summaries into claims for granular hallucination detection.
Like Vectara’s Hallucination Leaderboard Hughes and Bae (2023), other efforts like FACTS Grounding Jacovi et al. (2025) and Galileo’s Hallucination Index Galileo (2023) also provide leaderboards to benchmark hallucinations in LLMs. Galileo’s Hallucination Index employs GPT-4o as a single LLM judge, whereas FACTS Grounding ensembles evaluations from three different LLM judges: GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Pro.
@@
Nonetheless, hallucination detection remains challenging, with modest effectiveness observed across current methods. Benchmarks such as AggreFact, RAGTruth, TofuEval, and FaithBench consistently show limitations in existing hallucination detectors, including LLM-based methods. Notably, FaithBench highlighted that even the best current models achieve near 50% accuracy. Both RAGTruth and TofuEval further suggest that smaller, fine-tuned detection models can perform competitively with or even outperform LLM-based evaluation approaches.
@@
Table 1: Balanced Accuracy and F1-Macro of hallucination detection methods across four datasets. The final two columns report the simple average across the four datasets. We note that certain models marked with an asterisk (*) were used to select articles for the challenging FaithBench dataset.
3 Vectara’s Hallucination Leaderboard
In 2023, Vectara’s Hallucination Leaderboard Hughes and Bae (2023) was released using Vectara’s hallucination-detection model, HHEM-1.0-open. This model was later updated to HHEM-2.0 with stronger effectiveness, the ability to handle longer contexts, and multilingual capabilities. The current leaderboard relies on the open version, HHEM-2.1-open, publicly released on HuggingFace [^3] https://huggingface.co/vectara/hallucination_evaluation_model. To date, HHEM has been downloaded over 3.5 million times, reflecting strong community interest and adoption. While specific training details remain confidential, we note that HHEM-2.1-open was trained using the RAGTruth training set among other datasets.
To build Vectara’s Hallucination Leaderboard, articles were selected from diverse sources such as BBC News, CNN, Wikipedia, and the Daily Mail, following prior work on summarization evaluation and factuality verification Narayan et al. (2018); Maynez et al. (2020); Schuster et al. (2021); Thorne et al. (2018); Fabbri et al. (2021); Huang et al. (2020); Pagnoni et al. (2021); Hermann et al. (2015). Articles containing objectionable or explicit content, which LLMs may refuse to summarize, were specifically excluded. The resulting dataset comprised articles with a median length of approximately 217 words (25th percentile: 42 words; 75th percentile: 424 words).
LLMs are evaluated by prompting them to generate concise summaries strictly grounded in the provided passages. HHEM then assesses the proportion of summaries generated by the LLM containing hallucinations. Refusals are tracked by measuring the proportion of short responses (5 words or fewer). Users are also invited to submit specific models for evaluation. Continuously updated, the leaderboard now benchmarks hallucination rates of over 130 different LLMs, typically evaluating new models as soon as they become publicly available to track ongoing advances in the field.
4 FaithBench
FaithBench Bao et al. (2025) examined hallucinations in LLM-generated summaries and assessed the effectiveness of hallucination detection methods through human annotations. It includes summaries from ten state-of-the-art LLMs, including GPT-4o, GPT-3.5, Claude-3.5-Sonnet, Gemini-1.5-Flash Gemini Team (2024), and open-source models like Llama-3.1 Grattafiori et al. (2024), revealing that hallucinations remain frequent and detection methods often fail to identify them accurately.
Articles in FaithBench were selected from Vectara’s Hallucination Leaderboard based on frequent disagreements on summaries among hallucination detection models. True-NLI, TrueTeacher, HHEM-2.1-open, and GPT-4o/GPT-3.5 judges using the CoT prompt from Luo et al. (2023) were used to identify articles where summary hallucination classifications were most disagreed upon. The dataset includes 75 articles, each with ten annotated summaries from different LLMs.
@@ 4 FaithBench
   Majority Vote (Qwen 72B, Llama 70B, GPT-4o, o3) 80.2 81.3
Table 2: Balanced Accuracy and F1-Macro scores for FaithJudge on FaithBench using different LLM judges. With Majority Vote, we break ties by defaulting to a classification of inconsistent.
@@ 5 FaithJudge
Human annotation is the gold standard for hallucination detection, but it is time-consuming and expensive. FaithJudge offers a scalable alternative by leveraging hallucination annotations to guide an LLM judge in evaluating new summaries. We also expand FaithJudge to other RAG tasks, including question-answering (QA) and writing overviews from structured data in the JSON format using the RAGTruth dataset Niu et al. (2024). This is detailed further in the Appendix A.
To assess a summary, FaithJudge involves prompting an LLM judge with other summaries of the same article, along with their corresponding hallucination annotations. These annotations include hallucination spans, source references, and labels of either Benign, Unwanted, or Questionable, identified by human annotators.
@@ 6.1 Evaluation Datasets
We evaluate leading hallucination detection methods on four datasets: FaithBench, AggreFact Tang et al. (2023), RAGTruth Niu et al. (2024), and TofuEval-MeetingBank Tang et al. (2024b). While each of these datasets has previously analyzed hallucination detection individually, we provide a broader comparison across all four, motivating the need for our FaithJudge approach.
@@
For AggreFact, we evaluate on the SOTA subset of summaries, which involves annotated summaries generated by fine-tuned T5 Raffel et al. (2020), BART Lewis et al. (2020a), and PEGASUS Zhang et al. (2020) models. For RAGTruth, we evaluate only on the annotated summaries subset. Lastly, for TofuEval-Meetingbank, we evaluate on summaries generated using articles from the MeetingBank dataset Hu et al. (2023).
@@ 6.2 Existing Hallucination Detectors
Table 1 compares the effectiveness of fine-tuned hallucination detectors and zero-shot LLM-based methods across various datasets. We evaluate Vectara’s HHEM models alongside AlignScore, MiniCheck, including Bespoke-MiniCheck Bespoke (2024), and TrueTeacher. We also include current LLMs, such as GPT-4o and o3-mini (high reasoning), as well as open-source models Qwen2.5 (7B and 72B), Llama-3.1 (8B), and Llama-3.3 (70B). The o3-mini model, in particular, excels in reasoning tasks.
Classification methods are separated into claim-wise and summary-wise classification. Claim-wise evaluation involves decomposing sentences from summaries into individual claims using Llama-3.3 (70B) and a similar prompt from Tang et al. (2024a), while summary-wise methods assess the entire summary at once.
For LLM-based detection, we test three prompts: (1) the RAGAS prompt, which verifies lists of claims, (2) the FACTS Grounding JSON prompt, shown to be the most effective of the prompts tested in  Jacovi et al. (2025) for GPT-4o, and (3) the CoT-based prompt from Luo et al. (2023). We modify prompts slightly as needed for clearer final outputs and to specifically evaluate summaries.
@@
Unwanted $7474$ $322322$
Questionable $2929$ $3838$
Benign $5050$ $3434$
Consistent $176176$ $2727$
@@
Unwanted $8484$ $1818$ $294294$
Questionable $2828$ $1313$ $2626$
Benign $5151$ $1010$ $2323$
Consistent $179179$ $44$ $2020$
@@
Table 1 shows that, similar to findings in previous work, hallucination detection remains challenging. We note that certain models marked with an asterisk (*) were used to select articles for the adversarially challenging FaithBench dataset. Consequently, these models, including HHEM, may perform worse on FaithBench in summary-wise classification than they otherwise would. Zero-shot classification using GPT-4o and o3-mini-high tends to perform best, both using summary-wise classification with either the FACTS Grounding JSON prompt or the Luo et al. (2023) prompt. However, their average effectiveness remains modest, with balanced accuracy below 78% and F1-macro below 72%. Considering FaithBench, the highest balanced accuracy is achieved by o3-mini-high at 68.8% while the highest F1-macro of 63.7% is achieved by the HHEM model when considering claim-wise classification.
@@
Overall, fine-tuned models can score stronger than smaller prompted LLMs, but the largest LLMs typically yield the best results, even while being zero-shot methods. Regardless, the examined methods demonstrate modest effectiveness in general, with particularly weak effectiveness on FaithBench, which captures a diverse set of LLM summaries but is designed to be challenging for hallucination detection models.
Figure 2: Sensitivity and specificity with FaithJudge as the number of examples in the prompt are increased. We place an asterisk (*) next to the 10 because, in this case, FaithJudge is shown annotations for the summary it is evaluating.
Figure 3: Comparison of LLM rankings across FaithBench (based on (U) Unwanted, (B) Benign, and (Q) Questionable hallucination annotations), FaithJudge, and Vectara’s Hallucination Leaderboard. Rankings reflect the number of hallucinated summaries (from least to most).
6.3 FaithJudge
Table 2 presents the effectiveness of FaithJudge on FaithBench using various LLMs. The highest effectiveness is achieved using the o3-mini-high judge, reaching a balanced accuracy of 84% and an F1-macro of 82.1%, allowing for higher agreement with human annotation on FaithBench than the existing methods discussed in Table 1. Although the effectiveness of FaithJudge is not perfect, this may be partly explained by disagreements in human annotation. While human annotation is the gold standard, the FaithBench paper Bao et al. (2025) noted imperfect inter-annotator agreement in general and low inter-annotator agreement on more gray-area Benign and Questionable Hallucinations. We also note that we observe some erroneous annotations within FaithBench.
Effectiveness generally improves with increasing model size. We also tested an ensemble approach inspired by FACTS Grounding but found that combining predictions from multiple models, including o3-mini-high itself, did not outperform o3-mini-high alone. Therefore, we adopt the o3-mini-high judge as the standard for FaithJudge, with the possibility of using a stronger LLM judge down the line.

### v2

@@ Abstract
Retrieval-augmented generation (RAG) aims to reduce hallucinations by grounding responses in external context, yet large language models (LLMs) still frequently introduce unsupported information or contradictions even when provided with relevant context. This paper presents two complementary efforts at Vectara to measure and benchmark LLM faithfulness in RAG. First, we describe our original hallucination leaderboard, which has tracked hallucination rates for LLMs since 2023 using our HHEM hallucination detection model. Motivated by limitations observed in current hallucination detection methods, we introduce FaithJudge, an LLM-as-a-judge framework that leverages a pool of diverse human-annotated hallucination examples to substantially improve the automated hallucination evaluation of LLMs. We introduce an enhanced hallucination leaderboard centered on FaithJudge that benchmarks LLMs on RAG faithfulness in summarization, question-answering, and data-to-text generation tasks. FaithJudge enables a more reliable benchmarking of LLM hallucinations in RAG and supports the development of more trustworthy generative AI systems: https://github.com/vectara/FaithJudge.
@@
Manveer Singh Tamber1thanks: Work done while at Vectara., Forrest Sheng Bao3∗, Chenyu Xu2,3, Ge Luo2, Suleman Kazi2, Minseok Bae4∗, Miaoran Li3∗, Ofer Mendelevitch2, Renyi Qu2, Jimmy Lin1 1 University of Waterloo  2 Vectara  3 Iowa State University  4 Stanford University Correspondence: {mtamber, jimmylin}@uwaterloo.ca, ofer@vectara.com, forrest.bao@gmail.com
@@ 1 Introduction
Large language models (LLMs) excel in various tasks, but frequently produce hallucinations, generating false or misleading information unsupported by provided contexts or world knowledge hallucinationsurvey; hallucinationsurvey2; truthfulqa; aggrefact. While retrieval-augmented generation (RAG) approaches pmlr-v119-guu20a; rag; shuster-etal-2021-retrieval-augmentation seek to mitigate hallucinations by grounding responses in trusted contexts, they do not fully eliminate hallucinations, as LLMs often introduce details unsupported by retrieved contexts, misrepresent information, or generate outright contradictions ragtruth.
An ongoing challenge within RAG is evaluating and ensuring context-faithfulness ragtruth; jia-etal-2023-zero; ming2024faitheval. In this paper, we mainly focus on evaluating faithfulness in summarization tasks, building upon extensive prior research on summary consistency evaluation. Summarization tasks provide a practical benchmark for faithfulness, thanks to rich available hallucination datasets and established automated evaluation methods. However, despite recent progress, both fine-tuned detection models and LLM-as-a-judge techniques llmjudge; luo2023chatgpt; jacovi2025facts continue to struggle with accurately identifying hallucinations in LLM outputs.
We present two complementary efforts at Vectara for measuring and benchmarking LLM faithfulness in RAG. First, we describe our original hallucination leaderboard, which since 2023 has tracked hallucination rates for LLMs, currently ranking over 160 LLMs on hallucinations in summarization tasks. Second, motivated by the limitations of current hallucination detection approaches, we introduce FaithJudge, an LLM-as-a-judge framework that leverages a pool of diverse, human-annotated hallucination examples to improve automated faithfulness evaluation.
FaithJudge leverages labelled hallucination annotations from diverse LLM generations to automate the evaluation of LLMs on their propensity to hallucinate when summarizing the same articles or using the same articles to respond to queries. This approach results in notably higher agreement with human judgments compared with existing automated methods.
While our main investigation focuses on summarization tasks, we expand FaithJudge to other RAG tasks (including QA and data-to-text generation) using RAGTruth ragtruth, further detailed in Section 7.
Unlike previous work in the automated evaluation of LLMs for faithfulness in RAG, FaithJudge recasts LLM hallucination evaluation: the judge LLM studies a handful of human-annotated peer responses to the same source and then rules on a fresh candidate. Because the judge learns directly from the examples in its prompt, it needs no extra training, adapts naturally to new domains and tasks given annotations, and reaches state-of-the-art agreement with human annotators compared to existing methods while remaining fully automated.
@@ 2 Background
Numerous datasets have been developed for evaluating hallucinations in summarization tasks. Earlier datasets, such as SummaC laban-etal-2022-summac and AggreFact aggrefact, aggregated multiple resources and standardized labels and classification taxonomies. However, these primarily focused on summaries from pre-ChatGPT models like fine-tuned T5 t5, BART lewis-etal-2020-bart, and PEGASUS pegasus, potentially limiting their relevance to contemporary LLMs that may produce more nuanced and difficult-to-identify hallucinations.
@@
Recent benchmarks address this limitation by incorporating summaries generated by modern LLMs. TofuEval tang-etal-2024-tofueval provided hallucination labels on topic-focused dialogue summarization tasks with LLMs including GPT-3.5-Turbo, Vicuna vicuna2023 and WizardLM xu2023wizardlm. Similarly, HaluEval halueval included ChatGPT-generated hallucinations across summarization, question-answering (QA), and dialogue tasks, while RAGTruth ragtruth also annotated responses from models including GPT-3.5, GPT-4 gpt4, Llama-2 touvron2023llama, and Mistral jiang2023mistral7b. FaithBench bao-etal-2025-faithbench presented human annotations of challenging hallucinations in summaries from 10 modern LLMs from 8 different model families (detailed further in Section 4).
@@
Due to limited large-scale, human-annotated data for training hallucination detectors, early detection methods relied heavily on natural language inference (NLI) or question-answering (QA) systems fabbri-etal-2022-qafacteval. For instance, SummaC aggregated sentence-level NLI entailment scores between document-summary sentence pairs. AlignScore alignscore extended this by training detection models on multiple semantic alignment tasks evaluated at the chunk level. MiniCheck minicheck addressed data scarcity by synthesizing hallucinated examples using GPT-4 for model training.
@@
The strong zero-shot instruction-following capabilities of modern LLMs have also enabled LLM-as-a-judge methods llmjudge; luo2023chatgpt; jacovi2025facts; gao2023humanlikesummarizationevaluationchatgpt. Instead of evaluating entire generated summaries, approaches like FACTSCORE min-etal-2023-factscore and RAGAS es-etal-2024-ragas decompose summaries into claims for granular hallucination detection.
Like our original hallucination leaderboard, efforts such as FACTS Grounding jacovi2025facts and Galileo’s Hallucination Index GalileoAI_Hallucination_Index also provide leaderboards to benchmark hallucinations in LLMs, relying on LLM judges employed in a zero-shot manner to evaluate responses.
LLM judges have also been employed to evaluate various aspects of RAG, including the relevance of retrieved passages 10.1145/3626772.3657707, citation faithfulness liu-etal-2023-evaluating, and the factuality and completeness of responses 10.1145/3726302.3730090.
@@
Nonetheless, hallucination detection in RAG remains challenging, with weak effectiveness observed across current methods. Benchmarks such as AggreFact, RAGTruth, TofuEval, and FaithBench consistently show limitations in existing methods, including LLM-based ones. Notably, FaithBench highlighted that current methods, including using LLMs for classification, achieved near 50% accuracy, suggesting negligible ability to identify hallucinated responses. Further, both RAGTruth and TofuEval suggest that smaller, fine-tuned detection models can perform competitively with or even outperform zero-shot LLM-based evaluation approaches.
3 Vectara’s Original Hallucination Leaderboard
In 2023, Vectara’s hallucination leaderboard was released using Vectara’s hallucination detection model, HHEM-1.0-open. This model was later updated to HHEM-2.0 with improved effectiveness, the ability to handle longer contexts, and multilingual capabilities. The current leaderboard relies on the open version, HHEM-2.1-open, publicly released on HuggingFace. To date, HHEM has been downloaded over 4 million times, reflecting strong community interest and adoption. While specific training details remain confidential, we note that HHEM-2.1-open was trained using the RAGTruth training set among other datasets.
@@ 3 Vectara’s Original Hallucination Leaderboard
Table 1: Balanced Accuracy and F1-Macro of hallucination detection methods across four datasets. The final two columns report the simple average across the four datasets. We note that certain models marked with an asterisk (*) were used to select articles for the adversarially challenging FaithBench dataset. Consequently, these models, including HHEM, may perform worse on FaithBench in summary-wise classification than they otherwise would.
To build our original hallucination leaderboard, articles were selected from diverse sources such as BBC, CNN, Wikipedia, and the Daily Mail, following prior work on summarization evaluation and factuality verification xsum; xsumfactuality; schuster-etal-2021-get; fever; summeval; polytope; frank; dailymailcnn. Articles containing objectionable content, which LLMs may refuse to summarize, were excluded. The resulting dataset is composed of 955 articles with a median length of approximately 217 words.
LLMs are evaluated by prompting them to generate concise summaries strictly grounded on the provided passages. HHEM then assesses the proportion of summaries generated by the LLM containing hallucinations. Refusals are tracked by measuring the proportion of short responses (5 words or fewer). Users are also invited to submit specific models for evaluation. Continuously updated, the leaderboard now benchmarks hallucination rates of over 160 different LLMs, typically evaluating new models as soon as they become publicly available to track ongoing advances in LLMs.
@@
FACTS Grounding Prompt
   GPT-4o ? 65.9 56.2
   o3-mini-high ? 68.8 60.7
luo2023chatgpt Prompt
   GPT-4o ? 62.5 50.6
   o3-mini-high ? 63.3 49.8
@@
Table 2: Balanced Accuracy and F1-Macro for FaithJudge on FaithBench using different LLM judges. When combining models, we apply majority voting and break ties by defaulting to a classification of Inconsistent.
4 FaithBench
FaithBench bao-etal-2025-faithbench examined hallucinations in diverse LLM-generated summaries and assessed the effectiveness of hallucination detection methods through human annotations. It included summaries from ten state-of-the-art LLMs, including GPT-4o, GPT-3.5, Claude-3.5-Sonnet, Gemini-1.5-Flash team2024gemini, and open-source models like Llama-3.1 grattafiori2024llama, revealing that hallucinations remain frequent and detection methods generally fail to identify them reliably or accurately.
Articles in FaithBench were selected based on frequent disagreements on summaries among hallucination detection models. True-NLI, TrueTeacher, HHEM-2.1-open, and GPT-4o/GPT-3.5 judges using the chain-of-thought (CoT) prompt from luo2023chatgpt were used to identify articles where summary hallucination classifications were most disagreed upon. The dataset includes 75 articles, each with ten annotated summaries from different LLMs, allowing for many diverse LLM summaries to be studied per article. We show that this diversity in summaries also proves to be useful for FaithJudge when judging new summaries.
@@ 5 FaithJudge
Human annotation is the gold standard for hallucination detection, but it is time-consuming and expensive. FaithJudge offers a scalable alternative by leveraging hallucination annotations to guide an LLM judge in evaluating new summaries. We also expand FaithJudge to other RAG tasks, including question-answering (QA) and writing overviews from structured data in the JSON format using the RAGTruth dataset ragtruth. This is detailed further in Section 7.
To assess an LLM’s response, FaithJudge involves prompting an LLM judge with other responses to the same prompt, along with their corresponding hallucination annotations. These annotations include hallucination spans, source references from the context, and labels of either Benign, Unwanted, or Questionable, identified by multiple human annotators.
@@ 6.1 Evaluation Datasets
We evaluate hallucination detection methods on four summarization datasets: FaithBench, AggreFact aggrefact, RAGTruth ragtruth, and TofuEval-MeetingBank tang-etal-2024-tofueval. While each of these datasets has previously analyzed hallucination detection individually, we provide a comparison across all four, motivating the need for our FaithJudge approach.
@@
For AggreFact, we evaluate on the SOTA subset, which involves annotated summaries generated by fine-tuned T5 t5, BART lewis-etal-2020-bart, and PEGASUS pegasus models. For RAGTruth, we evaluate on the annotated summaries. Lastly, for TofuEval, we evaluate on summaries generated using articles from the MeetingBank dataset meetingbank.
@@ 6.2 Existing Hallucination Detectors
Table 1 compares the effectiveness of fine-tuned hallucination detectors and zero-shot LLM-based methods across various datasets. We evaluate Vectara’s HHEM models alongside AlignScore, MiniCheck, including Bespoke-MiniCheck tang2024bespokeminicheck, and TrueTeacher. We also include current LLMs, used in a zero-shot setting, such as GPT-4o and o3-mini-high, as well as open-source models Qwen2.5 (7B and 72B), Llama-3.1 (8B), and Llama-3.3 (70B). The o3-mini model, in particular, excels in reasoning tasks.
Classification methods are separated into claim-wise and summary-wise classification. Claim-wise evaluation involves decomposing sentences from summaries into individual claims using Llama-3.3 (70B) and a similar prompt from minicheck, while summary-wise methods assess the entire summary at once.
For LLM-based detection, we test three prompts: (1) the RAGAS prompt es-etal-2024-ragas, which verifies lists of claims instead of entire responses, (2) the FACTS Grounding JSON prompt, shown to be the most effective of the prompts tested in  jacovi2025facts for GPT-4o, and (3) the CoT-based prompt from luo2023chatgpt. We modify prompts slightly as needed for clearer final outputs and to specifically evaluate summaries.
Table 1 shows that, similar to findings in previous work, hallucination detection remains challenging. Zero-shot classification using GPT-4o and o3-mini-high tends to perform best, both using summary-wise classification with either the FACTS Grounding JSON prompt or the luo2023chatgpt prompt. However, their average effectiveness remains modest, with balanced accuracy below 78% and F1-macro below 72%. Considering FaithBench, the highest balanced accuracy is achieved by o3-mini-high at 68.8% while the highest F1-macro of 63.7% is achieved by the HHEM model when considering claim-wise classification.
@@
Unwanted $74$ $322$
Questionable $29$ $38$
Benign $50$ $34$
Consistent $176$ $27$
@@
Unwanted $84$ $18$ $294$
Questionable $28$ $13$ $26$
Benign $51$ $10$ $23$
Consistent $179$ 44 2020
@@
Overall, fine-tuned models can achieve stronger scores than smaller prompted LLMs, but the largest LLMs typically yield the best results, even while being zero-shot methods. This suggests room for further improvement with LLM-based classification by learning from hallucination labels. Regardless, the examined methods demonstrate modest effectiveness in general, with particularly weak effectiveness on FaithBench, which captures a diverse set of LLM summaries but is designed to be challenging for hallucination detection models.
Figure 2: Sensitivity and specificity with FaithJudge as the number of examples in the prompt is increased. We place an asterisk (*) next to the 10 because, in this case, FaithJudge is shown annotations for the summary it is evaluating.
6.3 Evaluating FaithJudge
Table 2 presents the effectiveness of FaithJudge on FaithBench using various LLMs. The highest effectiveness is achieved using the o3-mini-high judge, reaching a balanced accuracy of 84% and an F1-macro of 82.1%, allowing for much higher agreement with human annotation on FaithBench than the existing hallucination detection methods discussed. Although the effectiveness of FaithJudge is not perfect, this may be partly explained by disagreements in human annotation. While human annotation is the gold standard, the FaithBench authors  bao-etal-2025-faithbench noted imperfect inter-annotator agreement in general and low inter-annotator agreement on more ambiguous and challenging to identify Benign and Questionable hallucination labels.
Effectiveness generally improves with increasing model size. We also tested an ensemble approach, but found that combining predictions from multiple models, including Qwen2.5 (72B), Llama-3.3 (70B), and GPT-4o with a majority vote did not outperform o3-mini-high alone. Therefore, we adopt the o3-mini-high judge as the standard for FaithJudge, with the possibility of using a stronger LLM judge in the future.